OpenStudy (anonymous):

Use complete sentences to describe the steps taken to simplify this problem. Make sure you include information about the new common denominator, the final simplified expression, and any restrictions. 3/x + 3/x+1 - 3/x+2

OpenStudy (anonymous):

\[\frac{ 3 }{ x }+\frac{ 3 }{ x+1 }-\frac{ 3 }{ x+2 }\]